• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:01:55 on February28,2001

    Entry number 57178

    keyword=Shift Summary

    At start of shift had Hall access for polarity change and investigation of right Q3 He level trips. (False trips from incorrect level readings. Ad Hoc solution is to deliberately overfill)
    Spectrometers set for coincidence 15 kinematics
    Left:
    P0 = 2.638 (Positive polarity) at 29.971 deg
    Right:
    P0 = 2.291 (Negative polarity) at 35.193 deg.

    ep measurement attempted and abandoned by Joe Mitchell due to target failure.

    ARC integral measurement implies E = 4236.4 (see Halog entry 57160)

    Inspection of BeO shows spot 1mm left and 3mm high (see Halog 57159)

    run.........start/stop..........events..........comments
    2268......11:46/12:20.......0.5M.......LD2 with 6% radiator
    2269......12:26/13:09.......0.5M.......LD2 no radiator

    Change to LH2 target.
    Observations indicate problems with the H-loop fan. (see Halog 57173)

    2270.......13:49/ 14:24.......0.5M........LH2 no radiator
    2271.......14.24/15:15........1.0M.........LH2 with 6% radiator
    change targets
    2272.......15:27/15:28.........just a few.....LD2 with 6% radiator prescales incorrect
    2273.......15:30 handed off to swing shift